# This is the extended version of the conference version at FoSSaCS 2022 [BFPZ22] [a] [a] [a]

### v2

@@ 6.1.3. Finding quality optimal reachability PR causes
Now, we want to find recall- and coverage ratio-optimal reachability SPR causes. As in the state-based setting, we define the set $\mathcal{C}$ of all possible singleton reachability SPR causes for $\mathsf{rEff}$ that in addition satisfy (TempPrio) as explained in Remark 47. By Corollary 43, we can check whether a state $c\in S$ is a singleton reachability SPR cause in coNP; whether there is at least one state with $c$ in the first component in $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$ such that $\mathsf{noeff}_{\mathsf{fp}}$ is reachable from this state is checkable in polynomial time. Thus, we can again define the set of singleton reachability SPR causes $\mathcal{C}=\{s\in S\mid s\text{ is reachability SPR cause satisfying }\mathrm{Pr}^{\min}_{\mathcal{M}}(\mathsf{rEff}\mid\neg\mathsf{Cause}\,\mathrm{U}\,c)<1\}$. As before, the canonical cause $\mathsf{CanCause}$ is now the set of states $c\in\mathcal{C}$ for which there is a scheduler $\mathfrak{S}$ with $\mathrm{Pr}^{\mathfrak{S}}(\neg\mathcal{C}\,\mathrm{U}\,c)$.
@@ 6.2.1. Checking co-safety causality
$\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\mathsf{rCause}\land\mathsf{rEff})$ $=\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}_{[\mathsf{rEff},\mathsf{rCause}]}}(\lozenge\mathsf{eff}_{\mathsf{cov}})=\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}_{[\mathsf{rEff},\mathsf{rCause}]}}(\lozenge\mathsf{rCause}_{\mathcal{A}_{\mathsf{rEff}}}\land\lozenge E).$ (30)
Hence, $\mathsf{rCause}_{\mathcal{A}_{\mathsf{rEff}}}$ is a GPR cause for $E$ in $\mathcal{M}_{[\mathsf{rEff},\mathsf{rCause}]}$ if and only if $\mathsf{rCause}$ is a regular GPR cause for $\mathsf{rEff}$ in $\mathcal{M}$.
@@
$\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}_{[\mathsf{rEff},\mathsf{rCause}]}}(\lozenge E\mid\neg\mathsf{rCause}_{\mathcal{A}_{\mathsf{rEff}}}\,\mathrm{U}\,c)=\frac{\sum_{\pi\in\Pi}\sum_{\sigma\in\Sigma_{\pi}}\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\mathsf{rEff}\mid\sigma)\cdot\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\sigma)}{\sum_{\pi\in\Pi}\sum_{\sigma\in\Sigma_{\pi}}\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\sigma)}.$
All the terms $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\mathsf{rEff}\mid\sigma)$ are greater than $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\mathsf{rEff})=\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}_{[\mathsf{rEff},\mathsf{rCause}]}}(\lozenge E)$. Hence, we conclude that
@@
and thus $c$ is a state-based SPR cause for $\mathsf{eff}$. On the other hand for the scheduler $\tau$, which chooses $\alpha$ after the path $\pi=a\;b\;c$ and $\beta$ otherwise ,we have

### v6

@@ 6.1.3. Finding quality optimal reachability PR causes
Now, we want to find recall- and coverage ratio-optimal reachability SPR causes. As in the state-based setting, we define the set $\mathcal{C}$ of all possible singleton reachability SPR causes for $\mathsf{rEff}$ that in addition satisfy (TempPrio) as explained in Remark 47. By Corollary 43, we can check whether a state $c\in S$ is a singleton reachability SPR cause in coNP; whether there is at least one state with $c$ in the first component in $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$ such that $\mathsf{noeff}_{\mathsf{fp}}$ is reachable from this state is checkable in polynomial time. Thus, we can again define the set of singleton reachability SPR causes
$\mathcal{C}=\{s\in S\mid s\text{ is reachability SPR cause satisfying }\mathrm{Pr}^{\min}_{\mathcal{M}}(\mathsf{rEff}\mid\neg\mathsf{Cause}\,\mathrm{U}\,c)<1\}.$
As before, the canonical cause $\mathsf{CanCause}$ is now the set of states $c\in\mathcal{C}$ for which there is a scheduler $\mathfrak{S}$ with $\mathrm{Pr}^{\mathfrak{S}}(\neg\mathcal{C}\,\mathrm{U}\,c)$.
@@ 6.2.1. Checking co-safety causality
$\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\mathsf{rCause}\land\mathsf{rEff})$ $=\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}_{[\mathsf{rEff},\mathsf{rCause}]}}(\lozenge\mathsf{eff}_{\mathsf{cov}})=\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}_{[\mathsf{rEff},\mathsf{rCause}]}}(\lozenge\mathsf{rCause}_{\mathcal{A}_{\mathsf{rEff}}}\land\lozenge E),$ (30)
where $E=\{\mathsf{eff}_{\mathsf{cov}},\mathsf{eff}_{\mathsf{unc}}\}$. Hence, $\mathsf{rCause}_{\mathcal{A}_{\mathsf{rEff}}}$ is a GPR cause for $E$ in $\mathcal{M}_{[\mathsf{rEff},\mathsf{rCause}]}$ if and only if $\mathsf{rCause}$ is a regular GPR cause for $\mathsf{rEff}$ in $\mathcal{M}$.
@@
$\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}_{[\mathsf{rEff},\mathsf{rCause}]}}(\lozenge E\mid\neg\mathsf{rCause}_{\mathcal{A}_{\mathsf{rEff}}}\,\mathrm{U}\,c)=\frac{\sum_{\pi\in\Pi}\sum_{\sigma\in\Sigma_{\pi}}\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\mathsf{rEff}\mid\sigma)\cdot\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\sigma)}{\sum_{\pi\in\Pi}\sum_{\sigma\in\Sigma_{\pi}}\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\sigma)},$
where $E=\{\mathsf{eff}_{\mathsf{cov}},\mathsf{eff}_{\mathsf{unc}}\}$. All the terms $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\mathsf{rEff}\mid\sigma)$ are greater than the value of the term $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\mathsf{rEff})=\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}_{[\mathsf{rEff},\mathsf{rCause}]}}(\lozenge E)$. Hence, we conclude that
@@
and thus $c$ is a state-based SPR cause for $\mathsf{eff}$. On the other hand for the scheduler $\tau$, which chooses $\alpha$ after the path $\pi=\mathsf{init}\;b\;c$ and $\beta$ otherwise ,we have